Calculate the following four problems using interest tables, formulas, excel and the calculator. Thus, each problem (1 to 4) should be completed using four different methods. Please show all work for each of the four methods for each problem on the same excel sheet(s).

1) Determine the future value of $100,000 at 4% for 6 years

2) Determine the present value of $100,000 at 6% for 4 years

3) If you are to receive $7,419 each year (ordinary annuity) for the next 6 years, determine what this annuity is worth right now assuming 6% interest

4) What is the future value of $7,419 a year for 4 years at 6% interest
Answer #1

ANSWER

1)
Future value = $100000*FVIF @ 4% 6 years
= $100000*1.265 = $126,500
2)
Present value = $100000*PVIF @ 6% 4 years
= $100000*0.792 = $79,200
3)
Annuity is worth = $7,419 *FVIFA @6% 6 years
= $7,419 *6.975= $51,747.52
4)
Future value = $7,419 *FVIF @ 6% 4 years
= $7,419 *1.262 = $9,362.78

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************